Eight oil drops of same size are charged to a potential of 50 V each.These oil drops are merged into one single large drop.
What will be the potential of the large drop?Sir explain it with answer.

Grade:12th pass

3 Answers

Ravneet
5 years ago
Volume of 8 drop of radius ‘r’ = Volume of big drop of radius ‘R’ 
       8 X  4/3pi(r3 )= 4/3pi(R3)
            R = 2r 
potental of small drop  = kq/r = 50
potential of big drop = kq/2r = 50/2 = 25V
Ravneet
5 years ago
potential of bigger drop = k(8q)/2r = 4kq/r   Becouse R = 2r
and kq/r = 50 V 
so potential of bigger drop = 4 X 50  =  200 V
Chaitanya Prakhar Sharma
5 years ago
Use 
Potential =(n^2/3) kq/R=(n^2/3)v
n=no of drops
v=voltage given
put n=8 and v=50
Let's check its working
According to our problem
Potential of big drop = (8^2/3)*50
                                      =  4*50
                                      =200
Welcome!!!
 
 
 
 
 
 
 
 
 
 
Thank you...